Given the following tax structure, what minimum tax would need to be assessed on Shameika to make the tax progressive with respect to average tax rates?

Taxpayer Salary Muni-Bond Interest Total Tax
  Mihwah $22,500     $4,850          $1,665   
  Shameika $87,500     $40,000              ???   

 

Manimum tax ?
Answer #1
Mihwah Averge tax Rate = 1665/22500 *100 = 7.4%

So, Average TAx forShameika = 87500*7.4% = $6475

Thus, Shameika must pay more than 6475 (6476 or more) for the tax structure to be progressive for the average tax rates.